Operations Report – March 2nd

Crew 206 Operations Report 02 – MAR – 2019

SOL: 6

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 14,3 hours

Solar— SOC % – Turned off Gen : 69%, Turned on Gen : 54%

Diesel Reading – 75%

Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Gallons transferred: Not Applicable

Water in GreenHab – 257 gallons

Static to Loft Pump used – yes

Water Meter: 0141632,4 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s charger was down again (there might be an issue with water isolation somewhere as it happened on the two most wet days yet). Resetting the socket fixed it successfully.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A
ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Still on the fixing with BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Cleaned the whole Hab this morning. We tried to put as many recipients for water leaks during today’s "storm". Also, many tarps seem to have moved because of the wind. We plan on a full check and reparation on the next EVA.

The line of electric sockets supporting the water heater stopped working without triggering any circuit breaker. Resetting the socket fixed it seemingly.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing more to report.

Journalist Report – March 2nd

Crew 206 – 03/02/2019

Sol 6

Author: Auzou Benjamin, Journalist

"Telehope"

The return of the rain kept us inside the Hab for the morning. So we decided to clean both the lower and upper deck. The sand and the dust of the desert are our enemies, these particles are everywhere here and difficult to remove, but with courage, and a vigorous vacuum cleaner, we managed to make the Hab almost brand-new.

Cerise prepared us a bunch of cookies for the dessert. Even if she found that the cookies were "different from those she cook at home", they were delicious!

The afternoon marked the return of the hope for the experiment TELEOP. The digital version is still unavailable because of the loss of the computer’s charger, but Jérémy managed to correct the connection issues of the rover. We all drove the rover under heart monitoring.

This is the second time this experiment is performed here in the MDRS by ISAE-SUPAERO crews; it has also been developed by Simon Bouriat, Health and Safety Officer of MDRS 175 in the Lunares station (Poland) in 2018. This is a special experiment for us and our association. It’s pleasant to see the rover run again and we’re waiting for a charger to continue the digital experiment.

Tomorrow is our first day-off, no EVA, no experiments. We will use this free time to reorganize the Hab and realize cohesion activities, like playing cards or board games.

This is a tradition of ISAE-SUPAERO crews, like for the astronauts in the ISS, Sundays are off. We consider that this free time is indispensable to the success of such missions, psychology and human factors are important. Astronauts need this time for themselves and for crew cohesion.

Greenhab Report – March 2nd

Crew 206 Green Hab Report – 02-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (2 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.5°C
High: 26.1°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 5.3 gallons

Water in Blue Tank: 257.4 gallons

Crops watered at 0900 hours (1.3 gallons) and 1830 hours (4.0 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.00 AM 24.2° C / 27 %
2) 11.00 AM 25.4° C / 30 %
3) 15:00 PM 22.1° C / 45 %
4) 18:30 PM 23.8° C / 47 %

Changes to crops: We planted some new cress, with Aurélien and Benjamin, and also the tomatoes seeds.

Narrative:
Today, with the rain, I had to close the door at 1100, and to shut down the fan monitoring system.

As we couldn’t go in EVA, we planted two gardeners of cress with Aurélien and Benjamin.

In the afternoon, I planted the new seeds of tomatoes.

Harvest: None.

Support/supplies needed: None.

Sol Summary – March 2nd

Crew 206 March 2nd 2019
Sol 6

Summary Title : TELEHOPE

Author : Jeremy Auclair

Mission Status: All nominal

Sol Activity Summary: Waking up was even harder today. Even if we didn’t have a planned EVA we got up at 07:00 to workout. After breakfast, we spent a little more than an hour to clean the Hab, upper and lower deck. It was necessary, there was dust everywhere. Before lunch, half of us went to the GreenHab to plant some watercress, the rest of us helped Cerise to bake some cookies. It started raining when we started preparing lunch. After that most of us went to take a nap.

After 02:30 pm I started working on TELEOP again. I still don’t understand the connection issues I’m having, the experiment doesn’t work in the conditions I tested in Toulouse before leaving or even in the conditions that worked during MDRS mission 189. But I managed to find a set-up that worked: the rover and the subject in the science dome. It’s not ideal, it might change the analysis of the results but at least it allowed me to get some data. The simulated rover’s computer still doesn’t work, but we are not losing hope!

Tonight is going to be a rest night, we are going to watch a French comedy movie.

Look ahead plan: Tomorrow is our rest day, we will find some nice board games to play and sleep as much as we can.

Anomalies in work: None

Weather: Cloudy all day, started raining at 12h

Crew physical status: Fine

EVA : None, Bad Weather

Reports to be filed: GreenHab, Journalist, Operation

Support Requested: Birthday candles for two incoming birthdays (if it’s possible), Swiffer dust removers.